OpenStudy (anonymous):

Find all values of x that make the equation true. | = absolute value a. | 3x+5 |=1 b. | 6-x |=10

OpenStudy (lgbasallote):

| 3x + 5 | = 1 solve for x in 3x + 5 = 1 and 3x + 5 = -1 does that help?

OpenStudy (anonymous):

yes, thank you!

OpenStudy (anonymous):

so is it simply -(4/3) and -2?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

same thing for number 2 solve for x in 6 - x = 10 and 6- x = -10

OpenStudy (anonymous):

yup, got it